THE MILLIMETERWAVE SPECTROSCOPY OF THE VINYL-2-d_{1} RADICAL (HDC=CH)

MASATO HAYASHI, KENSUKE HARADA and \underline{KEIICHI TANAKA}, Department of Chemistry, Faculty of Sciences, Kyushu University, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

The vinyl-2-d_{1} radical HDC=CH, mono-deuteride of the methylenic (CH_{2}) proton, may have two forms, cis- and trans-HDCCH; for cis-HDCCH an unpaired electron is located on the same side of deuterium with respect to the C=C double bond, while on the opposite side for trans-HDCCH. If the barrier hight h between the cis- and trans-form is sufficiently low and the zero point energy differece between the two forms Δ_{tc0} is small enough (Δ_{tc0} ≪ 1 cm^{−1}), the proton on the acetylenic side (CH) will easily transfer through the barrier by tunneling effect, giving us a chance to observe the tunneling transitions between the two forms (0^{+} and 0^{−}). If not, the two forms will remain as two toutomers, cis-HDCCH and trans-HDCCH, and will give us a chance to observe the b-type rotational transitions within the cis and trans toutomers instead of the tunneling transitions. In our previous study ^{a}, we have observed the a-type rotational spectra of cis-HDCCH by millimeter wave spectroscopy in a supersonic jet, while no lines for trans-HDCCH were observed. In the present study, we have searched for either the tunneling transitions or the b-type rotational transitions of HDCCH generated by 193 nm excimer laser photolysis of HDCCHCl. The b-type rotational transitions (N_{K_{a}K_{c}}=1_{11}−0_{00}, 2_{12}−1_{01}, 1_{10}−1_{01} and 2_{11}−2_{02}) only for one form of HDCCH were observed in the frequency region of 158−258 GHz, but no the tunneling transitions. The rotational and hyperfine interction constants, ^{b} indicate that the observed toutomer is not trans-HDCCH but cis-HDCCH.

Our ab initio calulation performed in the CCSD(T)/cc-pVQZ level supports the present observation. Althogh the calculated barrier hight between the cis- and trans-form is as low as h = 1770 cm^{−1}w, the zero point energy gap between the two toutomers Δ_{tc0} = 30.7 cm^{−1} is large, as the result we have two toutomers, cis-HDCCH and trans-HDCCH, to observe the b-type rotational transitions instead of the tunneling transitions. The zero point energy gap Δ_{tc0} 30.7 cm^{−1} is large enough to populate only to the vibrational ground state of cis-HDCCH in the ultra low temperature T = 15K in the supersonic jet giving us no trans-HDCCH spectra.

aM. Hayashi, K. Harada, K. Tanaka, International Symposium on Molecular Spectroscopy, TJ14(2006)

bK. Tanaka, M. Toshimitsu, K. Harada, T. Tanaka, J. Chem. Phys.120, 3604 (2004)